OpenStudy (anonymous):

How would you find a slope on a line Im talking about a X and Y axis coordinate with say (-1,-1 to 2,2) how would I find the slope what does it mean?

OpenStudy (jtvatsim):

slope = rise/run or more formally, slope = (y2 - y1)/(x2 - x1) For example, given (1, 2) and (3, 4) slope = (4 - 2)/(3 - 1) = 2/2 = 1